# assumeAssumption

### v1

@@ Abstract
What does it mean to be flat? We propose to define it by measuring the maximal variation around a point, or from a dual perspective, the distance to neighboring level sets. After developing some calculus rules, we show how flat minima, conservation laws, and symmetries are intertwined. Gradient flows of conserved quantities are of particular interest, due to their flattening properties. Keywords: dynamical systems, semi-algebraic geometry, variational analysis.
@@ 1 Introduction
Flat minima were informally introduced by Hochreiter and Schmidhuber [19, 20] in the context of deep learning as connected regions of minima where the objective is nearly constant. In a desire to formalize this idea, it was later suggested to use the volume of connected components of sublevel sets [13, Definition 1]. Around the same time, an empirical measure of sharpness [26, Metric 2.1] was proposed to analyze the role of batch size in the stochastic subgradient method for training neural networks. For nonnegative functions, it was redefined as the maximal ratio of the function variation over one plus the function value, over a ball of fixed radius [13, Definition 2].
@@
where $\mathrm{ReLU}(t)=\max\{0,t\}$. The objective function is differentiable at its global minima, so its gradient is defined and equal to zero (details in Example˜6). But it is not twice differentiable there, so its Hessian is not defined. As for the volumes of sublevel sets, they are generally not finite on ReLU networks [13, Theorem 2], nor on the most simple neural network, i.e., $f(x)=(x_{1}x_{2}-1)^{2}$.
@@
obeys $\nabla f(x_{1},0)=(0,0)$ and $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$ for all $x_{1}\in\mathbb{R}$, according to which all the global minima $(x_{1},0)$ are allegedly flat. But factoring in the fourth-order growth actually suggests that $(0,0)$ is the unique flat minimum (see Example˜8). This is not merely a theoretical matter, as gradient descent with sufficiently slowly diminishing step lengths converges to the origin. This can be seen in Fig.˜1, and is proved in the sequel [22].
@@
This clarifies the picture in matrix factorization
@@
where $M\in\mathbb{R}^{m\times n}$, $\|\cdot\|_{F}$ is the Frobenius norm, and $C(X,Y)=X^{T}X=YY^{T}$. Some authors [12] report the flat minima are “balanced”, meaning that $X^{T}X=YY^{T}$, while others [33] say that they are “nearly balanced”. With our definition, we find that while $X^{T}X-YY^{T}$ can be nonzero, there does exist a balanced maximal eigenvector of the Hessian $\nabla^{2}f(X,Y)$. In passing, we note that flat minima admit a simple characterization. Namely, if $XY=M$, then
@@
where $\|\cdot\|_{2}$ is the spectral norm. Our analysis also reveals a new local-global property of matrix factorization. While it is known the every local minimum is a global minimum [5, 38], it was not known that locally flat minima are globally flat.
Third, we determine flat minima in a series of examples that were out of reach with previous definitions and tools. They demonstrate that the definition of flatness proposed in this work is both versatile and workable. We also illustrate the connection with symmetry and conservation. In one example,
$f(x)=(x_{1}^{\upsilon_{1}}\cdots x_{n}^{\upsilon_{n}}-1)^{2}$
where $\upsilon_{1},\ldots,\upsilon_{n}$ are positive integers, any gradient trajectory of the conserved quantity initialized at a global minimum converges to a flat global minimum. It does so while remaining in the (n1)(n-1)-dimensional solution manifold.
@@ 2 Background
Given a matrix $M\in\mathbb{R}^{m\times n}$, $M^{T}$ denotes the transpose. If $A:U\to V$ is a linear map between two finite dimensional inner product spaces, then $A^{*}:V\to U$ denotes the adjoint. Let $\|\cdot\|_{F}$, $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, $\|\cdot\|_{*}$, $\|\cdot\|_{p,q}$, and $\rho(\cdot)$ respectively denote the Frobenius norm, entrywise $\ell_{1}$-norm, spectral norm, nuclear norm, $(p,q)$-induced norm, and spectral radius. Also, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius norm. Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, $\lambda_{1}(M)$ and $E_{1}(M)$ respectively denote its maximal eigenvalue and its associated eigenspace, whose nonzero elements are referred to as maximal eigenvectors.
@@ 2.2 Differential calculus
Given some vector spaces $V_{1},\ldots,V_{k}$ and $W$, a map $F:V_{1}\times\cdots\times V_{k}\rightarrow W$ is multilinear if it is linear in each variable taken separately when the others are held fixed. It is symmetric if $V_{1}=\cdots=V_{k}$ and $F(x_{\sigma(1)},\ldots,x_{\sigma(k)})=F(x_{1},\ldots,x_{k})$ for any permutation $\sigma$ of $\llbracket 1,k\rrbracket$. When it is real-valued, i.e., $W=\mathbb{R}$, $F$ is called a multilinear form. Given a symmetric multilinear map $F:V\times\cdots\times V\to W$ and norms $\|\cdot\|_{V},\|\cdot\|_{W}$ on $V,W$, consider the norm [7, Theorem A]
@@
where $o:\mathbb{R}_{+}\to W$ is a function such that $o(t)/t\to 0$ at $t\searrow 0$. We say that $f$ is $C^{1}$ at $\overline{x}$ if $f$ is $D^{1}$ near $\overline{x}$ and $f^{\prime}$ is continuous at $\overline{x}$. By induction, for any integer $k\geq 2$, we may define $f$ to be $D^{k}$ (resp. $C^{k}$) at $\overline{x}$ if it is $D^{k-1}$ (resp. $C^{k-1}$) near $\overline{x}$ and $f^{(k-1)}$ is $D^{1}$ (resp. $C^{1}$) at $\overline{x}$. We only define $D^{k}$ and $C^{k}$ for $k\in\mathbb{N}^{*}$. We say $f$ is $D^{k}$ (resp. $C^{k}$) if it is so at every point in $U$. Also, $f$ is $C^{k,\ell}$ (resp. $C^{k,\ell}_{L}$) if it is $C^{k}$ and $f^{(\ell)}$ is locally Lipschitz continuous (resp. $L$-Lipschitz continuous). If $f$ is $D^{k}$ at $\overline{x}$, then $f^{(k)}$ is a symmetric multilinear map [17, Proposition C.16]. When evaluated on the diagonal, we use the shorthand $f^{(k)}(x)v^{k}=f^{(k)}(x)(v,\ldots,v)$.
@@ 2.3 Differential geometry
An action of a group $G$ with identity $e$ on a set $M$ [27, p. 161] is a map $\theta:G\times M\rightarrow M$ such that
@@
When an action exists, $M$ is referred to as a $G$-space. It is homogeneous if for all $x,y\in M$, there exists $g\in G$ such that $\theta(g,x)=y$. A function $f:M\rightarrow N$ between sets $M$ and $N$ is invariant under an action of a group $G$ on $M$ if $f(gx)=f(x)$ for all $g\in G$ and $x\in M$.
A Lie group $G$ is a smooth manifold and a group whose operations are smooth. A Lie group $G$ acts smoothly on a smooth manifold $M$ if there exists a smooth action $\theta:G\times M\rightarrow M$. A Lie subgroup of a Lie group $G$ is a subgroup of $G$ is endowed with a topology and smooth structure making into a Lie group and an immersed submanifold. Topologically closed subgroups of Lie groups are Lie subgroups by the closed subgroup theorem [27, Theorem 20.12]. Let $\mathfrak{g}$ denote the Lie algebra of a Lie group $G$, which we identify with its tangent space at $e$.
Let $I_{n}$ denote the identity matrix of order $n$. The set of invertible matrices with real coefficients of order $n$, denoted $\mathrm{GL}(n,\mathbb{R})$, is a Lie group. The natural action of a Lie subgroup $G$ of $\mathrm{GL}(n,\mathbb{R})$ on $\mathbb{R}^{n}$ is defined by the matrix vector multiplication $G\times\mathbb{R}^{n}\ni(g,x)\mapsto gx\in\mathbb{R}^{n}$. The orthogonal group $\mathrm{O}(n)=\{Q\in\mathbb{R}^{n\times n}:Q^{T}Q=I_{n}\}$ is a Lie subgroup of $\mathrm{GL}(n,\mathbb{R})$.
@@ 2.4 Variational analysis
It will be convenient to use the generalization of local optiamlity from points to sets proposed in [25, 24]. A nonempty set $S\subseteq\mathbb{R}^{n}$ is a local minimum (resp. strict) of $f:\mathbb{R}^{n}\to\mathbb{R}$ if there exists a neighborhood $U$ of $S$ such that $f(x)\leq f(y)$ (resp. $f(x)<f(y)$) for all $x\in S$ and $y\in U\setminus S$. In constrast to [24], we do not assume $S$ to be closed.
@@
and define other expressions like $f\leq\ell]$ similarly. Let
$\mathrm{dom}f=\{x\in\mathbb{R}^{n}:f(x)<\infty\}~~~\text{and}~~~\mathrm{gph}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)=t\}.$
The indicator of $S\subseteq\mathbb{R}^{n}$ is defined by
$\delta_{S}(x)=\left\{\begin{array}[]{cl}0&\text{if}~x\in S,\\ \infty&\text{if}~x\notin S.\end{array}\right$.
A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is lower semicontinuous at $\overline{x}\in\mathrm{dom}f$ if $\liminf_{x\rightarrow\overline{x}}f(x)\geq f(\overline{x})$ [36, Definition 1.5]. It is lower semicontinuous if it is so at every point in its domain. The regular normal cone and normal cone [36, Definition 6.3] are defined by
@@
A set $C\subseteq\mathbb{R}^{n}$ is regular [36, Definition 6.4] at one of its points $\overline{x}$ if it is locally closed and $\widehat{N}_{C}(\overline{x})=N_{C}(\overline{x})$.
Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the regular subdifferential, subdifferential [36, Definition 8.3], and Clarke subdifferential of $f$ at $\overline{x}$ [14, Definition 4.1] are
@@
where $\mathrm{co}$} denotes the convex hull, and $\overline{\mathrm{co}}$ its closure. The $o$ means that $\liminf[f(x)-f(\overline{x})-\langle v,x-\overline{x}\rangle]/|x-\overline{x}|\geq 0$ where $\overline{x}\neq x\to\overline{x}$. A point $x\in\mathbb{R}^{n}$ is critical (resp. Clarke critical) if $0\in\partial f(x)$ (resp. $0\in\overline{\partial}f(x)$). A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is regular [36, Definition 7.25] at $\overline{x}$ if $f(\overline{x})$ is finite and $\mathrm{epi}f$ is regular at $(\overline{x},f(\overline{x}))$ as a subset of $\mathbb{R}^{n+1}$. The Lipschitz modulus of a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined by [36, p. 354]
@@
where $D$ are the differentiable points of $f$. In what amounts to a slight abuse of notation, we may replace $D$ by some $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}$ has zero Lebesgue measure. Still assuming $\mathrm{lip}f(\overline{x})<\infty$, by [36, Theorem 9.13] $\partial f(\overline{x})$ is nonempty and compact, and $\mathrm{lip}f(\overline{x})=\max\{|v|:v\in\partial f(\overline{x})\}$. By [36, Theorem 9.61], $\overline{\nabla}f(\overline{x})$ is a nonempty compact subset of $\partial f(\overline{x})$ and $\mathrm{co}\overline{\nabla}f(\overline{x})=\mathrm{co}\hskip 0.85358pt\partial f(\overline{x})=\overline{\partial}f(\overline{x})$. Thus $\arg\max\{|v|:v\in\partial f(\overline{x})\}=\arg\max\{|v|:v\in\overline{\nabla}f(\overline{x})\}$, using the following fact.
Fact 1.

### v2

# On the geometry of flat minima

@@ Abstract
What does it mean to be flat? We propose to define it by measuring the maximal variation around a point, or from a dual perspective, the distance to neighboring level sets. After developing some calculus rules, we show how flat minima, conservation laws, and symmetries are intertwined. Gradient flows of conserved quantities are of particular interest, due to their flattening properties.
Keywords: dynamical systems, semi-algebraic geometry, variational analysis.
@@ 1 Introduction
Flat minima were informally introduced by Hochreiter and Schmidhuber [22, 23] in the context of deep learning as connected regions of minima where the objective is nearly constant. In a desire to formalize this idea, it was later suggested to use the volume of connected components of sublevel sets [15, Definition 1]. Around the same time, an empirical measure of sharpness [31, Metric 2.1] was proposed to analyze the role of batch size in the stochastic subgradient method for training neural networks. For nonnegative functions, it was redefined as the maximal ratio of the function variation over one plus the function value, over a ball of fixed radius [15, Definition 2].
@@
where $\mathrm{ReLU}(t)=\max\{0,t\}$. The objective function is differentiable at its global minima, so its gradient is defined and equal to zero (details in Example˜5.1). But it is not twice differentiable there, so its Hessian is not defined. As for the volumes of sublevel sets, they are generally not finite on ReLU networks [15, Theorem 2], nor on the most simple neural network, i.e., $f(x)=(x_{1}x_{2}-1)^{2}$.
@@
obeys $\nabla f(x_{1},0)=(0,0)$ and $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$ for all $x_{1}\in\mathbb{R}$, according to which all the global minima $(x_{1},0)$ are allegedly flat. But factoring in the fourth-order growth actually suggests that $(0,0)$ is the unique flat minimum (see Example˜5.3). This is not merely a theoretical matter, as gradient descent with sufficiently slowly diminishing step lengths converges to the origin. This can be seen in Fig.˜1, and is proved in the sequel [27].
@@
Third, we determine flat minima in a series of examples that were out of reach with previous definitions and tools. They demonstrate that the definition of flatness proposed in this work is both versatile and workable. We also illustrate the connection with symmetry and conservation.
This work clarifies the picture in overparamatrized matrix factorization
@@
where $M\in\mathbb{R}^{m\times n}$, $\|\cdot\|_{F}$ is the Frobenius norm, $r\geq\mathrm{rank}\hskip 0.85358pt(M)$, and $C(X,Y)=X^{T}X=YY^{T}$. Some authors [14] report the flat minima are “balanced”, meaning that $X^{T}X=YY^{T}$, while others [38] say that they are “nearly balanced”. With our definition, we find that balanced global minima are flat, but the converse is false. While $X^{T}X-YY^{T}$ can be nonzero at flat minima, there exists a balanced maximal eigenvector of the Hessian $\nabla^{2}f(X,Y)$. In passing, we note that flat minima admit a simple characterization. Namely, if $XY=M$, then
@@
where $\|\cdot\|_{2}$ is the spectral norm.
Our analysis also reveals a new local-global property of matrix factorization. While it is known the every local minimum is a global minimum [5, 46], it was not known that locally flat minima are globally flat. A related novelty is that any gradient trajectory of the conserved quantity $\|C\|_{F}^{2}$ initialized at a global minimum converges to a flat global minimum.
@@ 2 Background
Given a matrix $M\in\mathbb{R}^{m\times n}$, $M^{T}$ denotes the transpose. If $A:U\to V$ is a linear map between two finite dimensional inner product spaces, then $A^{*}:V\to U$ denotes the adjoint. Let $\|\cdot\|_{F}$, $\|\cdot\|_{1}$, $\|\cdot\|_{2}$, $\|\cdot\|_{*}$, $\|\cdot\|_{p,q}$, and $\rho(\cdot)$ respectively denote the Frobenius norm, entrywise $\ell_{1}$-norm, spectral norm, nuclear norm, $(p,q)$-induced norm, and spectral radius. Also, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius norm. Given a symmetric matrix $M\in\mathbb{R}^{n\times n}$, $\lambda_{1}(M)$ and $E_{1}(M)$ respectively denote its maximal eigenvalue and its associated eigenspace, whose nonzero elements are referred to as maximal eigenvectors. We use $M\succcurlyeq 0$ to denote that $M$ is positive semidefinite.
@@ 2.1 Ordered sets
Let $S$ be a subset of an ordered set $X$. An element $y\in X$ is a least (resp. greatest) element of $S$ if $y\in S$ and $x\leq y$ (resp. $y\leq x$) for all $x\in S$. A lower (resp. upper) bound of $S$ is an element $y\in X$ such that $y\leq x$ (resp. $x\leq y$) for all $x\in S$. A lower bound of $S$ is an infimum of $S$ if it is a greatest lower bound. If an infimum exists, then it is unique. An infimum is a minimum if it belongs to $S$. The set of infima and minima are respectively denoted $\inf S$ and $\min S$. Suprema and maxima are defined analogously, and denoted $\sup S$ and $\max S$ respectively.
A totally ordered set is complete if every nonempty subset that has a lower bound, has a greatest lower bound. The real numbers $\mathbb{R}$ is such an example. Let $S\subseteq\mathbb{R}$. If $S$ has no lower bound, then $\inf S=$-$\infty$. If $S=\emptyset$, then $\inf S=\infty$.
@@ 2.2 Differential calculus
Given some vector spaces $V_{1},\ldots,V_{k}$ and $W$, a map $F:V_{1}\times\cdots\times V_{k}\rightarrow W$ is multilinear if it is linear in each variable taken separately when the others are held fixed. It is symmetric if $V_{1}=\cdots=V_{k}$ and $F(x_{\sigma(1)},\ldots,x_{\sigma(k)})=F(x_{1},\ldots,x_{k})$ for any permutation $\sigma$ of $\llbracket 1,k\rrbracket$. When it is real-valued, i.e., $W=\mathbb{R}$, $F$ is called a multilinear form. Given a symmetric multilinear map $F:V\times\cdots\times V\to W$ and norms $\|\cdot\|_{V},\|\cdot\|_{W}$ on $V,W$, consider the norm [8, Theorem A]
@@
where $o:\mathbb{R}_{+}\to W$ is a function such that $o(t)/t\to 0$ at $t\searrow 0$. We say that $f$ is $C^{1}$ at $\overline{x}$ if $f$ is $D^{1}$ near $\overline{x}$ and $f^{\prime}$ is continuous at $\overline{x}$. By induction, for any integer $k\geq 2$, we may define $f$ to be $D^{k}$ (resp. $C^{k}$) at $\overline{x}$ if it is $D^{k-1}$ (resp. $C^{k-1}$) near $\overline{x}$ and $f^{(k-1)}$ is $D^{1}$ (resp. $C^{1}$) at $\overline{x}$. We only define $D^{k}$ and $C^{k}$ for $k\in\mathbb{N}^{*}$. We say $f$ is $D^{k}$ (resp. $C^{k}$) if it is so at every point in $U$. Also, $f$ is $C^{k,\ell}$ (resp. $C^{k,\ell}_{L}$) if it is $C^{k}$ and $f^{(\ell)}$ is locally Lipschitz continuous (resp. $L$-Lipschitz continuous). If $f$ is $D^{k}$ at $\overline{x}$, then $f^{(k)}$ is a symmetric multilinear map [19, Proposition C.16]. When evaluated on the diagonal, we use the shorthand $f^{(k)}(x)v^{k}=f^{(k)}(x)(v,\ldots,v)$.
@@ 2.3 Differential geometry
An action of a group $G$ with identity $e$ on a set $M$ [32, p. 161] is a map $\theta:G\times M\rightarrow M$ such that
@@
When an action exists, $M$ is referred to as a $G$-space. It is homogeneous if for all $x,y\in M$, there exists $g\in G$ such that $\theta(g,x)=y$. A function $f:M\rightarrow N$ between sets $M$ and $N$ is invariant under an action of a group $G$ on $M$ if $f(\theta(g,x))=f(x)$ for all $g\in G$ and $x\in M$.
A Lie group $G$ is a smooth manifold and a group whose operations are smooth. A Lie group $G$ acts smoothly on a smooth manifold $M$ if there exists a smooth action $\theta:G\times M\rightarrow M$. A Lie subgroup of a Lie group $G$ is a subgroup of $G$ is endowed with a topology and smooth structure making into a Lie group and an immersed submanifold. Topologically closed subgroups of Lie groups are Lie subgroups by the closed subgroup theorem [32, Theorem 20.12]. Let $\mathfrak{g}$ denote the Lie algebra of a Lie group $G$, which we identify with its tangent space at $e$.
Let $I_{n}$ denote the identity matrix of order $n$. The set of invertible matrices with real coefficients of order $n$, denoted $\mathrm{GL}(n)$, is a Lie group. The natural action of a Lie subgroup $G$ of $\mathrm{GL}(n)$ on $\mathbb{R}^{n}$ is defined by the matrix vector multiplication $G\times\mathbb{R}^{n}\ni(g,x)\mapsto gx\in\mathbb{R}^{n}$. The orthogonal group $\mathrm{O}(n)=\{Q\in\mathbb{R}^{n\times n}:Q^{T}Q=I_{n}\}$ is a Lie subgroup of $\mathrm{GL}(n)$.
@@ 2.4 Variational analysis
It will be convenient to use the generalization of local optimality from points to sets proposed in [29, 28]. A nonempty set $S\subseteq\mathbb{R}^{n}$ is a local minimum (resp. strict) of $f:\mathbb{R}^{n}\to\mathbb{R}$ if there exists a neighborhood $U$ of $S$ such that $f(x)\leq f(y)$ (resp. $f(x)<f(y)$) for all $x\in S$ and $y\in U\setminus S$. In constrast to [28], we do not assume $S$ to be closed.
@@
and define other expressions like $f\leq\ell]$ similarly. Let $\mathrm{dom}f=\{x\in\mathbb{R}^{n}:f(x)<\infty\}$,
$\mathrm{gph}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)=t\},~~~\text{and}~~~\mathrm{epi}f=\{(x,t)\in\mathbb{R}^{n}\times\mathbb{R}:f(x)\leq t\}.$
The indicator and the support function of a set $C\subseteq\mathbb{R}^{n}$ are respectively defined by
$\forall x\in\mathbb{R}^{n},~\delta_{C}(x)=\left\{\begin{array}[]{cl}0&\text{if}~x\in C,\\ \infty&\text{if}~x\notin C.\end{array}\right.~~~\text{and}~~~\forall v\in\mathbb{R}^{n},~\sigma_{C}(v)=\sup_{w\in C}\langle v,w\rangle.$
Given $C\subseteq\mathbb{R}^{n}$ and $\overline{x}\in C$, the tangent cone [44, Definition 6.1] is defined by
$T_{C}(\overline{x})=\{v\in\mathbb{R}^{n}:~\exists x_{k}\xrightarrow[C]{}\overline{x}~,~\tau_{k}\searrow 0:(x^{k}-\overline{x})/\tau^{k}\to v\}$
where $x_{k}\xrightarrow[C]{}\overline{x}$ is a shorthand for $x_{k}\rightarrow\overline{x}$ and $x_{k}\in C$. A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is lower semicontinuous at $\overline{x}\in\mathrm{dom}f$ if $\liminf_{x\rightarrow\overline{x}}f(x)\geq f(\overline{x})$ [44, Definition 1.5]. It is lower semicontinuous if it is so at every point in its domain. The regular normal cone and normal cone [44, Definition 6.3] are defined by
@@
A set $C\subseteq\mathbb{R}^{n}$ is regular [44, Definition 6.4] at one of its points $\overline{x}$ if it is locally closed and $\widehat{N}_{C}(\overline{x})=N_{C}(\overline{x})$. The polar set of $C\subseteq\mathbb{R}^{n}$ [6, Definition 6.22] is defined by
$C^{*}=\{v\in\mathbb{R}^{n}:~\forall w\in C,~\langle v,w\rangle\leq 0\}.$
By [44, Theorem 6.28], the relationship $\widehat{N}_{C}(\overline{x})=T_{C}(\overline{x})^{*}$ holds.
Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the one-sided directional derivative (if it exists in $\mathbb{R}\cup\{\pm\infty\}$) and the subderivative for $w\in\mathbb{R}^{n}$ are respectively given by
$f^{\prime}(\overline{x};w)=\lim_{\tau\searrow 0}\frac{f(\overline{x}+\tau w)-f(\overline{x})}{\tau}~~~\text{and}~~~df(\overline{x})(w)=\liminf_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\overline{x}+\tau w^{\prime})-f(\overline{x})}{\tau}.$
According to [44, Definition 7.20], $f$ is semidifferential at $\overline{x}$ for $w\in\mathbb{R}^{n}$ if
$\lim_{\scriptsize\begin{array}[]{c}\tau\searrow 0\\ w^{\prime}\to w\end{array}}\frac{f(\overline{x}+\tau w^{\prime})-f(\overline{x})}{\tau}\in\mathbb{R}\cup\{\pm\infty\}.$
If this holds for all $w\in\mathbb{R}^{n}$, then $f$ is semidifferentiable at $\overline{x}$. Given $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ and a point $\overline{x}\in\mathbb{R}^{n}$ where $f(\overline{x})$ is finite, the regular subdifferential, subdifferential, horizon subdifferential [44, Definition 8.3], and Clarke subdifferential of $f$ at $\overline{x}$ [16, Definition 4.1] are respectively given by
@@
$\partial^{\infty}f(\overline{x})=\{v\in\mathbb{R}^{n}:\exists(x_{k},v_{k})\in\mathrm{gph}\hskip 1.42262pt\widehat{\partial}f:\exists\tau_{k}\searrow 0:(x_{k},f(x_{k}),\tau_{k}v_{k})\rightarrow(\overline{x},f(\overline{x}),v)\},$
@@
where $\mathrm{co}$} denotes the convex hull, and $\overline{\mathrm{co}}$ its closure. The $o$ means that
$\liminf_{\scriptsize\begin{array}[]{c}x\rightarrow\overline{x}\\ x\neq\overline{x}\end{array}}\frac{f(x)-f(\overline{x})-\langle v,x-\overline{x}\rangle}{|x-\overline{x}|}\geq 0.$
A point $x\in\mathbb{R}^{n}$ is critical (resp. Clarke critical) if $0\in\partial f(x)$ (resp. $0\in\overline{\partial}f(x)$). A function $f:\mathbb{R}^{n}\rightarrow\overline{\mathbb{R}}$ is regular [44, Definition 7.25] at $\overline{x}$ if $f(\overline{x})$ is finite and $\mathrm{epi}f$ is regular at $(\overline{x},f(\overline{x}))$ as a subset of $\mathbb{R}^{n+1}$. The Lipschitz modulus of a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is defined by [44, p. 354]
@@
where $D$ are the differentiable points of $f$. In what amounts to a slight abuse of notation, we may replace $D$ by some $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}$ has zero Lebesgue measure. Still assuming $\mathrm{lip}f(\overline{x})<\infty$, by [44, Theorem 9.13] $\partial f(\overline{x})$ is nonempty and compact, and $\mathrm{lip}f(\overline{x})=\max\{|v|:v\in\partial f(\overline{x})\}$. By [44, Theorem 9.61], $\overline{\nabla}f(\overline{x})$ is a nonempty compact subset of $\partial f(\overline{x})$ and $\mathrm{co}\overline{\nabla}f(\overline{x})=\mathrm{co}\hskip 0.85358pt\partial f(\overline{x})=\overline{\partial}f(\overline{x})$. Thus $\arg\max\{|v|:v\in\partial f(\overline{x})\}=\arg\max\{|v|:v\in\overline{\nabla}f(\overline{x})\}$, using the following fact.
Fact 2.1.